OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–1, 3) and (1, 7)? y – 7 = 4(x – 1) y – 7 = 2(x – 1) y – 3 = 2(x – 1) y – 3 = 4(x + 1)

OpenStudy (blurbendy):

you need to find the slope first. slope = (y2 - y1) / (x2 - x1) where y2 = 7 y1 = 3 x2 = 1 x1 = -1

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

oh no its 2

OpenStudy (blurbendy):

7 -3 = 4 1 + 1 = 2 slope = 2, yes

hero (hero):

So eliminate two of your answer choices.

OpenStudy (blurbendy):

now use the formula: y - y1 = 4(x - x1) use either (-1, 3) or (1,7) for x1 and y1, but do not combine them

hero (hero):

Its between: y – 7 = 2(x – 1) or y – 3 = 2(x – 1)

OpenStudy (blurbendy):

that 4 should be a 2, sorry

hero (hero):

@blurbendy, I thought we agreed the slope was 2

OpenStudy (anonymous):

is it y-7 = 2(x-1)?

OpenStudy (blurbendy):

yes, now isolate y

OpenStudy (anonymous):

thats all i need, thank you!

OpenStudy (blurbendy):

actually, you dont need to. youre done!